# Fractal and Regular Geometry of Deep Neural Networks

### v1

@@ A.3   Technical lemmas
$A_{k,q;n}^{(s)}=(-1)^{(k-1)/2+q}\frac{2^{k}n!}{k!q!(n-k-q)!}\ ,$ $k\text{ odd }\ ,$
$A_{k,q;n}^{(s)}=0\ ,$ $k\text{ even }\ ,$
$A_{k,q;n}^{(c)}=(-1)^{k/2+q}\frac{2^{k}n!}{k!q!(n-k-q)!}\ ,$ $k\text{ even }\ ,$
$A_{k,q;n}^{(c)}=0\ ,$ $k\text{ odd }\,$
@@
[…]
@@
  • Suppose $d$ is odd, with $d=2n+1$. From the previous lemma and noting that, for $20<\psi<\theta<\pi/$ 2, we have
@@
where $Q_{0,a;d}$ and $Q_{1,a;d}$ are the same constants given in Lemma A.5.
@@
    hence, using the definition of $A_{k,q;n}^{[i]}$ the claim follows.
@@
Proposition 4.4 is a generalization to any arbitrary dimension of Theorem 1 in [37]. To obtain the results in any dimension, one can follow the proof in [37] using Proposition B.3 above instead than their Proposition 7. To prove Proposition B.3, let us recall that every sequence of spherical harmonics satisfies the following addition formula (cfr. [6], Theorem 2.9)
@@
The next lemma ensures that such sequence exists.
@@
which proves the claim. For the inductive step, note that following the construction in [6] Section 2.11, if $(Y_{\ell,m;d-1})$ is an orthonormal basis of real spherical harmonics in dimension $d-1$, then we can construct a basis of real spherical harmonics in $\mathbb{S}^{d}$: $(Y_{\ell,m;d})$, where
@@
[…]
@@
[…]
@@
Using the addition formula (cf. (B.1)), we obtain
@@
The next lemma is instrumental to the proof of Proposition B.3.
@@
Let $T$ be an isotropic random field with $\beta<1$ as CRI. For all $10<\varepsilon<$ 1 there exists a triangular sequence $(\kappa_{\ell,m;d}(\varepsilon)\;|\;\ell\in\mathbb{N},\ m=1,\dotsc,n_{\ell,d})$ such that
@@
    $=\frac{1}{\varepsilon^{d}(d-1)!\omega_{d}}\sum_{\ell=1}^{\infty}G(x_{\ell})x_{\ell}^{d-1}(x_{\ell+1}-x_{\ell})\cdot\left(\frac{\ell}{\ell+d-1}\right)^{(d-1)/2}\frac{\sqrt{(\ell+1)(\ell+d)}+\sqrt{\ell(\ell+d-1)}}{2\ell+d}$
@@
    Since $(x_{\ell})_{\ell=L}^{\infty}$ forms a partition of $[\varepsilon\sqrt{L(L+d-1)},\infty)$, we have
@@
    From Huygens’ principle ([30], Lemma 4.1) if $K_{\varepsilon}$ is the kernel of the previous operator, then
@@
Under the assumptions of Proposition 4.4, we have

### v2

@@ A.3 Technical lemmas
[…]
@@
$Q_{2,a;d}$ $=-2^{(d-5)/2+a}\operatorname{B}\left(\frac{a+3}{2},\frac{d-1}{2}\right)\frac{3d^{2}-9d+3-a}{3}\quad$ $\text{ $d$ even}\ ,$
$Q_{2,a;d}$ $=-2^{(d-5)/2+a}\operatorname{B}\left(\frac{a+3}{2},\frac{d-1}{2}\right)\frac{3d^{2}-9d-a}{3}\quad$ $\text{ $d$ odd}\ .$
@@
  • Suppose $d$ is odd, with $d=2n+1$. From the previous lemma and noting that, for $0<\psi<\theta<\pi/2$, we have
@@
where $Q_{0,a;d}$ and $Q_{1,a;d}$ are the same constants given in Lemma˜A.5.
@@
    hence, using the definition of $A_{k,q;n}^{[i]}$, the claim follows..
@@
Proposition˜4.5 is a generalization to any arbitrary dimension of Theorem 2 in [40]. To obtain the results in any dimension, one can follow the proof in [40] using Proposition˜B.3 above instead of their Proposition 7. To prove Proposition˜B.3, let us recall that every sequence of spherical harmonics satisfies the following addition formula (see [8], Theorem 2.9)
@@
The next lemma ensures that such a sequence exists.
@@
which proves the claim. For the inductive step, note that following the construction in [8] Section 2.11, if $(Y_{\ell,m;d-1})$ is an orthonormal basis of real spherical harmonics in dimension d1d-1, then we can construct a basis of real spherical harmonics in $\mathbb{S}^{d}$: $(Y_{\ell,m;d})$, where
@@
$\qquad\qquad\qquad\qquad\qquad\vdots$
@@
$\qquad\qquad\qquad\qquad\qquad\vdots$
@@
Using the addition formula (cfr. (B.1)), we obtain
@@
The next lemma is instrumental to the proof of Proposition˜B.3.
@@
Let $T$ be an isotropic random field with $\beta<1$ as CRI. For all $0<\varepsilon<1$, there exists a triangular sequence $(\kappa_{\ell,m;d}(\varepsilon)\;|\;\ell\in\mathbb{N},\ m=1,\dotsc,n_{\ell,d})$ such that
@@
    $=\frac{1}{\varepsilon^{d}(d-1)!\omega_{d}}\sum_{\ell=1}^{\infty}G(x_{\ell})x_{\ell}^{d-1}(x_{\ell+1}-x_{\ell})\left(\frac{\ell}{\ell+d-1}\right)^{(d-1)/2}\frac{\sqrt{(\ell+1)(\ell+d)}+\sqrt{\ell(\ell+d-1)}}{2\ell+d}$
@@
    Since $(x_{\ell})_{\ell=L}^{\infty}$ forms a partition of $\varepsilon\sqrt{L(L+d-1)},\infty)$, we have
@@
    From Huygens’ principle [33, Lemma 4.1], if $K_{\varepsilon}$ is the kernel of the previous operator, then
@@
Under the assumptions of Proposition˜4.5, we have
@@
Appendix C Lower bound of (3.3): Hausdorff dimension of the excursion set
Let $\mathcal{M}^{+}$ be the set of all positive measures on $\mathbb{R}^{d+1}$. For any $\alpha>0$, we define the following inner product on $\mathcal{M}^{+}$,
$\langle\mu,\nu\rangle_{\alpha}=\iint_{(\mathbb{R}^{d+1})^{2}}\frac{1}{\|x-y\|^{d+1-\alpha}}\mathrm{d}\mu(x)\mathrm{d}\nu(y)\ ,$
and set $\mathcal{E}_{\alpha}^{+}=\{\mu\in\mathcal{M}^{+}\;|\;\|\mu\|_{\alpha}<\infty\}$. It is well-known that $(\mathcal{E}^{+}_{\alpha},\|\cdot\|_{\alpha})$ is a Hilbert space [41, Theorem 1.18]. The following proposition implies the lower bound of (3.3) as explained in Section˜4.5.
Proposition C.1.
Let $T$ be a zero mean, unit-variance isotropic Gaussian random field on $\mathbb{S}^{d}$. If the CRI $\beta\in(0,1]$, then, for every $u\in\mathbb{R}$, there exists a random measure $\mu$ on $\mathbb{S}^{d}$ such that, for every $\gamma<d-\beta$,
$\int_{(\mathbb{S}^{d})^{2}}\frac{1}{\|x-y\|^{\gamma}}\mathrm{d}\mu(x)\mathrm{d}\mu(y)<\infty\qquad\text{a.s.}$
and such that $\mu(T^{-1}(u))>0$ with positive probability.
Proof.
For any $n\in\mathbb{N}$, let $\mu_{n}$ be the random measure on $\mathbb{R}^{d+1}$ such that, for any Borel set $A\subseteq\mathbb{R}^{d+1}$,
$\mu_{n}(A)=\int_{A\cap\mathbb{S}^{d}}\phi_{n}(T(x))\mathrm{d}x,\qquad\text{where}\quad\phi_{n}(t)=\sqrt{{2n\pi}}e^{-\frac{n(t-u)^{2}}{2}}\,.$
Let $n,m\in\mathbb{N}$ and $\alpha>0$. Then, setting $\gamma=d+1-\alpha$ we have
$\mathbb{E}[\langle\mu_{n},\mu_{m}\rangle_{\alpha}]=\mathbb{E}\bigg[\int_{(\mathbb{R}^{d+1})^{2}}\frac{1}{\|x-y\|^{\gamma}}\mathrm{d}\mu_{n}(x)\mathrm{d}\mu_{m}(y)\bigg]$
$\qquad=2\pi\sqrt{nm}\mathbb{E}\bigg[\int_{(\mathbb{S}^{d})^{2}}\exp\Big(-\frac{1}{2}(n(u-T(x))^{2}+m(u-T(y))^{2})\Big)\|x-y\|^{-\gamma}\mathrm{d}x\mathrm{d}y\bigg]$
$\qquad=2\pi\sqrt{nm}\int_{(\mathbb{S}^{d})^{2}}\daleth(x,y)\|x-y\|^{-\gamma}\mathrm{d}x\mathrm{d}y\ ,$
where the last identity follows from Fubini’s theorem and
$\daleth(x,y)=\mathbb{E}\bigg[\exp\Big(-\frac{1}{2}(n(u-T(x))^{2}+m(u-T(y))^{2})\Big)\bigg]\ .$
Since $T$ is a zero-mean Gaussian random field with covariance kernel $\kappa$, then putting $\rho=\kappa(\langle x,y\rangle)$ we have
$(u-T(x),u-T(y))\sim\mathcal{N}(\bm{u},\Sigma),\qquad\bm{u}=\begin{pmatrix}u\\ u\end{pmatrix},\quad\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\$.
Hence, for $\bm{s}=(s_{1},s_{2})^{\top}\in\mathbb{R}^{2}$ we have
$\daleth(x,y)=$ $\int_{\mathbb{R}^{2}}\frac{1}{2\pi\sqrt{\det\Sigma}}e^{-\frac{1}{2}\big(ns_{1}^{2}+ms_{2}^{2}+(\bm{s}-\bm{u})^{\top}\Sigma^{-1}(\bm{s}-\bm{u})\big)}\mathrm{d}\bm{s}\ .$
Using some algebraic manipulation, one can prove that
$ns_{1}^{2}+ms_{2}^{2}+(\bm{s}-\bm{u})^{\top}\Sigma^{-1}(\bm{s}-\bm{u})=(\bm{s}-\bm{\mu}_{nm})^{\top}\Sigma_{mn}^{-1}(\bm{s}-\bm{\mu}_{nm})+h_{nm}$
where
$\Sigma_{mn}^{-1}=\frac{1}{1-\rho^{2}}\begin{pmatrix}n(1-\rho^{2})+1&-\rho\\ -\rho&m(1-\rho^{2})\end{pmatrix},\qquad\bm{\mu_{nm}}=\begin{pmatrix}\mu_{1}^{nm}\\ \mu_{2}^{nm}\end{pmatrix}$
$h_{nm}=u^{2}\frac{2mn(1-\rho)+m+n}{mn(1-\rho^{2})+m+n}$
with
$\mu_{1}^{nm}=u\frac{mn\rho^{2}-mn-m\rho-n}{mn\rho^{2}-mn-mn-m-n-1},\qquad\mu_{2}^{nm}=u\frac{mn\rho^{2}-mn-n\rho-m}{mn\rho^{2}-mn-m-n-1}\ .$
Hence
$\daleth(x,y)=\sqrt{\frac{\det\Sigma_{mn}}{\det\Sigma}}e^{-h_{nm}/2}=\frac{e^{-h_{nm}/2}}{\sqrt{nm(1-\rho^{2})+n+m+1}}\ .$
With this computation, we obtain
$\mathbb{E}[\langle\mu_{n},\mu_{m}\rangle_{\alpha}]$ $=2\pi\sqrt{nm}e^{-h_{nm}/2}\int_{(\mathbb{S}^{d})^{2}}\frac{\|x-y\|^{\alpha-d-1}}{\sqrt{nm(1-\rho^{2})+n+m+1}}\mathrm{d}x\mathrm{d}y$ (C.1)
$=Ce^{-h_{nm}/2}\int_{-1}^{1}\frac{(1-r)^{(\alpha-d-1)/2}}{\sqrt{1-\kappa(r)^{2}+\frac{1}{n}+\frac{1}{m}+\frac{1}{nm}}}(1-r^{2})^{d/2-1}\mathrm{d}r$
$\leq C\int_{-1}^{1}\frac{(1-r)^{(\alpha-3)/2}(1+r)^{d/2-1}}{(1-\kappa(r)^{2})^{1/2}}\mathrm{d}r\ .$
where $C$ is a constant independent of $n$ and $m$ and in the last inequality we use $h_{nm}\geq 0$. Schoenberg’s theorem implies that $|\kappa(r)|=1$ is possible only for $r\in\{-1,1\}$, so the previous integral may diverge only for $r$ near $\pm 1$.
  • Using the definition of CRI, as $r\to 1^{-}$, we have
    $1-\kappa(r)=c_{1}(1-r)^{\beta}+o((1-r)^{\beta})\ ,$
    and thus
    $\frac{(1-r)^{(\alpha-3)/2}(1+r)^{d/2-1}}{(1-\kappa(r)^{2})^{1/2}}\sim C(1-r)^{-1+(\alpha-1-\beta)/2}\ .$
    Hence the integral (near $1$) is finite if and only if $\alpha-1-\beta>0$, i.e., $\gamma=d+1-\alpha<d-\beta$.
  • For $r$ near 1-1, we consider three different cases:
    • If $\kappa(-1)\neq\pm 1$, then the integral (near 1-1) is finite.
    • If $\kappa(-1)=1$, then as $r\to(-1)^{+}$ we have
      $1-\kappa(r)=c_{-1}(1+r)^{\beta}+o((1+r)^{\beta})\ ,$
      and thus
      $\frac{(1-r)^{(\alpha-3)/2}(1+r)^{d/2-1}}{(1-\kappa(r)^{2})^{1/2}}\sim C(1+r)^{-1+(d-\beta)/2}\ .$
      Hence, since $\beta\in(0,1]$ and $d\geq 2$, the integral (near 1-1) is finite.
    • If $\kappa(-1)=-1$, then as $r\to(-1)^{+}$ we have
      $1+\kappa(r)=c_{-1}(1+r)^{\beta}+o((1+r)^{\beta})\ ,$
      and we obtain the same conclusion as in the previous case.
In particular, we see that taking $n=m$ and $\alpha>\beta+1$, we have $\mathbb{E}[\|\mu_{n}\|^{2}_{\alpha}]<\infty$. Therefore $\|\mu_{n}\|_{\alpha}<\infty$ almost surely, that is $\mu_{n}\in\mathcal{E}_{\alpha}^{+}$ a.s.. Let us now prove that the sequence $(\mu_{n})_{n}$ is a.s. Cauchy. Using (C.1),
$\mathbb{E}[\|\mu_{n}-\mu_{m}\|_{\alpha}^{2}]=C\int_{-1}^{1}(1-r)^{(\alpha-3)/2}(1+r)^{d/2-1}f_{nm}(r)\mathrm{d}r\ ,$
where
$f_{nm}(r)=$ $\frac{e^{-h_{nn}/2}}{\sqrt{1-\kappa(r)^{2}+\frac{2}{n}+\frac{1}{n^{2}}}}+\frac{e^{-h_{mm}/2}}{\sqrt{1-\kappa(r)^{2}+\frac{2}{m}+\frac{1}{m^{2}}}}-\frac{2e^{-h_{nm}/2}}{\sqrt{1-\kappa(r)^{2}+\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}}}\ .$
We note that $f_{nm}(r)\to 0$ as $n,m\to+\infty$. Moreover,
$|f_{nm}(r)|\leq\frac{4}{\sqrt{1-\kappa(r)^{2}}}\ .$
Since we have proved that
$\int_{-1}^{1}\frac{(1-r)^{(\alpha-3)/2}(1+r)^{d/2-1}}{\sqrt{1-\kappa(r)^{2}}}\mathrm{d}r<\infty\ ,$
using the dominated convergence theorem we get $\lim_{n,m\to\infty}\mathbb{E}[\|\mu_{n}-\mu_{m}\|_{\alpha}]=0$, and hence $(\mu_{n})_{n}$ is a Cauchy sequence in $\mathcal{E}_{\alpha}^{+}$ on an event $\Omega_{0}$ of full probability. Let $\mu$ be the limiting (random) measure. We claim that there exists a set $B\subseteq\Omega_{0}$ with positive probability such that $\mu(T^{-1}(u))>0$ on $B$.
First we prove that $\mu_{n}(\mathbb{S}^{d})>0$ with positive probability. To do this, we use the Paley–Zygmund inequality, i.e., for any random variable $X\geq 0$ and for any $\theta\in[0,1]$,
$\mathbb{P}(X>\theta\mathbb{E}[X])\geq(1-\theta)^{2}\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[X^{2}]}\ .$ (C.2)
Now,
$\mathbb{E}[\mu_{n}(\mathbb{S}^{d})]$ $=\mathbb{E}\Big[\int_{\mathbb{S}^{d}}\sqrt{2\pi n}e^{-\frac{n(u-T(x))^{2}}{2}}\mathrm{d}x\Big]=\int_{\mathbb{S}^{d}}\sqrt{2\pi n}\int_{\mathbb{R}}\frac{1}{\sqrt{2\pi}}e^{-\frac{n(u-w)^{2}}{2}}e^{-\frac{w^{2}}{2}}\mathrm{d}w\mathrm{d}x$
$=\sqrt{n}\omega_{d}e^{-\frac{n}{2(n+1)}u^{2}}\int_{\mathbb{R}}e^{-\frac{n+1}{2}\big(w-\frac{n}{n+1}u\big)^{2}}\mathrm{d}w$
$=\sqrt{2\pi}\omega_{d}e^{-\frac{n}{2(n+1)}u^{2}}\sqrt{\frac{n}{n+1}}\ .$
Since $\mu_{n}\in\mathcal{E}_{d+1}^{+}$, we also have $\mathbb{E}[\mu_{n}(\mathbb{S}^{d})^{2}]\leq C$. Using (C.2), we obtain
$\mathbb{P}(\mu_{n}(\mathbb{S}^{d})>\theta\mathbb{E}[\mu_{n}(\mathbb{S}^{d})])\geq\frac{(1-\theta)^{2}}{C}2\pi\omega_{d}^{2}e^{-\frac{n}{2(n+1)}u^{2}}\frac{n}{n+1}\geq p$
where $p$ is a positive constant independent of $n$. On the other hand,
$\{\mu_{n}(\mathbb{S}^{d})>\theta\mathbb{E}[\mu_{n}(\mathbb{S}^{d})]\}=\Big\{\mu_{n}(\mathbb{S}^{d})>\theta\sqrt{2\pi}\omega_{d}e^{-\frac{n}{2(n+1)}u^{2}}\sqrt{\frac{n}{n+1}}\Big\}\subseteq\{\mu_{n}(\mathbb{S}^{d})>q\},$
for $q>0$ independent of $n$. Thus, there exists $B\subseteq\Omega_{0}$ such that $\mathbb{P}(B)\geq p$ and $\mu_{n}(\mathbb{S}^{d})>q$ on $B$.
Finally, let us prove that $\mu(T^{-1}(u))>0$ with positive probability. We will even prove that $\mu$ has support in $T^{-1}(u)$ almost surely. Let $U$ be an open neighbourhood of $T^{-1}(u)$. Then $\mathbb{S}^{d}\setminus U$ is a closed set, and there exists
$\delta=\min_{t\in\mathbb{S}^{d}\setminus U}|T(t)-u|>0\,.$
Thus
$\phi_{n}(T(x))\leq\sqrt{2\pi n}e^{-\frac{n\delta^{2}}{2}}\leq\sqrt{2\pi}\delta^{-1}\mathrm{e}^{-1/2}\qquad\forall x\in\mathbb{S}^{d}\setminus U\,.$
On the other hand, $\phi_{n}(t)\to 0$ for any $t\neq u$, and hence, by dominated convergence,
$\mu(\mathbb{S}^{d}\setminus U)=\lim_{n\to+\infty}\mu_{n}(\mathbb{S}^{d}\setminus U)=\lim_{n\to+\infty}\int_{\mathbb{S}^{d}\setminus U}\phi_{n}(T(x))\mathrm{d}\mu_{n}(x)=0\ ,$
which concludes the proof. ∎